7. Shannon Company segments its income statement into its North and South Divisions. The company’s overall sales, contribution margin ratio, and net operating income are $480,000, 36%, and $9,600, respectively. The North Division’s contribution margin and contribution margin ratio are $92,000 and 40%, respectively. The South Division’s segment margin is $32,500. The company has $62,400 of common fixed expenses that cannot be traced to either division.

Required: Prepare an income statement for Shannon Company that uses the contribution format and is segmented by divisions. In addition, for the company as a whole and for each segment, show each item on the segmented income statements as a percent of sales. (Round your percentage answers to 1 decimal place (i.e .1234 should be entered as 12.3).
Answer #1

Answer

  • Requirement as asked

Total company

North

South

Amount

%

Amount

%

Amount

%

Sales

$480,000

100.0%

$230,000

100.0%

$250,000

100.0%

Variable expenses

$307,200

64.0%

$138,000

60.0%

$169,200

67.7%

Contribution margin

$172,800

36.0%

$92,000

40.0%

$80,800

32.3%

Traceable Fixed expenses

$100,800

21.0%

$52,500

22.8%

$48,300

19.3%

Territorial segment margin

$72,000

15.0%

$39,500

17.2%

$32,500

13.0%

Common Fixed expenses

$62,400

Net Operating Income

$9,600

--Working

Total company

North

South

Amount

Amount

Amount

%

Sales

480000

=92000/40%

=480000-230000

=+G37666/$G$37666

Variable expenses

=480000-172800

=230000-92000

=307200-138000

=+G37667/$G$37666

Contribution margin

=480000*36%

92000

=172800-92000

=+G37668/$G$37666

Traceable Fixed expenses

=172800-72000

=92000-39500

=100800-52500

=+G37669/$G$37666

Territorial segment margin

=62400+9600

=72000-32500

32500

=+G37670/$G$37666

Common Fixed expenses

62400

Net Operating Income

9600